# WEATHER-5K: A Large-scale Global Station Weather Dataset Towards Comprehensive Time-series Forecasting Benchmark

### v1

@@ Abstract
Global Station Weather Forecasting (GSWF) is crucial for various sectors, including aviation, agriculture, energy, and disaster preparedness. Recent advancements in deep learning have significantly improved the accuracy of weather predictions by optimizing models based on public meteorological data. However, existing public datasets for GSWF optimization and benchmarking still suffer from significant limitations, such as small sizes, limited temporal coverage, and a lack of comprehensive variables. These shortcomings prevent them from effectively reflecting the benchmarks of current forecasting methods and fail to support the real needs of operational weather forecasting. To address these challenges, we present the WEATHER-5K dataset. This dataset comprises a comprehensive collection of data from 5,672 weather stations worldwide, spanning a 10-year period with one-hour intervals. It includes multiple crucial weather elements, providing a more reliable and interpretable resource for forecasting. Furthermore, our WEATHER-5K dataset can serve as a benchmark for comprehensively evaluating existing well-known forecasting models, extending beyond GSWF methods to support future time-series research challenges and opportunities. The dataset and benchmark implementation are publicly available at: https://github.com/taohan10200/WEATHER-5K.
@@ 1 Introduction
Global Station Weather Forecasting (GSWF) is essential for providing precise and timely weather information, with significant implications for various sectors. Accurate weather forecasts enable airlines to enhance aviation safety [1], support agriculture in optimizing crop management [2], and assist the energy industry in managing production and mitigating weather-related risks [3]. Moreover, reliable forecasts are crucial for early warning systems, aiding in the preparation for natural disasters and extreme weather events, thereby safeguarding lives and property [4, 5].
However, the current GSWF predominantly relies on Numerical Weather Prediction (NWP) models, which include both physical-based NWP models [6, 7] and data-driven NWP models [8, 9, 10, 11]. These models do not optimally consider forecasting as an end-to-end optimized task, leading to two primary challenges. First, such models are computationally demanding, relying on extensive resources for data assimilation and medium-range forecasting. Second, the results obtained from these models are typically represented as grid data, which lacks the precision required to accurately represent specific weather station locations.
Recently, there have been some initial efforts [12] to formulate GSWF as an end-to-end task by directly learning from global station weather observations. However due to limitations of the public global station weather datasets, the scale of stations and data used for optimizing such methods remains relatively small. Specifically, current datasets often collect data from a single station [13] or a localized region [14] or limited time range [12]. This characteristic results in forecasting methods achieving only a one-day lead time, limiting their applicability in real-world scenarios. Hence, it is urgent to develop a comprehensive global station weather dataset that enables forecasting models to generalize across diverse stations and regions worldwide. We believe such a comprehensive dataset would significantly enhance the robustness and applicability of GSWF models, addressing the limitations of current approaches.
Moreover, such a large weather station dataset can also serve as an extensive time-series dataset to perform comprehensive time-series forecasting benchmarks for various forecasting methods proposed for diverse purposes (e.g., traffic prediction models [15, 16]). In fact, due to the lack of large-scale time-series datasets, most existing forecasting methods have primarily been evaluated and analyzed using small-scale datasets. These simple datasets fail to encompass the complex scientific problems that researchers need to discover and resolve, thereby hindering progress in the field of time-series prediction. Our dataset offers a diverse temporal and spatial range of time-series data, enabling a comprehensive evaluation of time-series forecasting methods and driving significant advancements in the field. To summarize, this research offers two significant contributions, namely:
@@
  • Introduction of the WEATHER-5K Dataset. We propose a large-scale time-series dataset for sparse weather forecasting, namely WEATHER-5K, which comprises 5,672 weather stations located worldwide. This extensive dataset provides a diverse representation of weather conditions across different regions. It also offers substantial temporal coverage, with each station containing 10 years of hourly data, enabling the analysis of long-term weather patterns and the development of forecasting models that capture seasonal and interannual variations. We also conduct thorough data analysis to uncover trends, patterns, and correlations within our weather data.
  • Extensive Time-Series Benchmarking. We implement a comprehensive set of widely recognized time-series forecasting methods encompassing diverse domains such as traffic prediction, weather prediction, and electricity prediction. To evaluate the performance of different forecasting algorithms, we conducted extensive benchmark experiments on WEATHER-5K, establishing a standardized evaluation framework. The empirical findings derived from these benchmarks not only shed light on research challenges but also identify future opportunities for the advancement of GSWF and time-series forecasting.
In the future, our newly proposed dataset can support operational weather services for the public, an essential aspect of weather forecasting research that often holds greater significance than the research itself. Additionally, weather station data serve as a crucial source of observational data for NWP models, effectively bridging the gap between numerical models and station-based predictions. This not only improves the accuracy of numerical forecasts but also plays a vital role in verifying and evaluating the predictive performance of NWP models [11].
Figure 1: : An illustration of the WEATHER-5K benchmark dataset.
@@ 2 Related Work
2.1 Time-series Forecasting
@@
Time-series forecasting involves many domains [14, 17, 18, 19] and their methods have undergone three stages since its birth: statistical learning, machine learning, and deep learning.
@@
Statistical learning methods, including ARIMA [20], ETS [21], StatsForecast [22], VAR [14], and Kalman Filter (KF) [23], are among the early proposals and are widely utilized. These methods rely on historical data to predict future values and are based on the assumption that past observations hold predictive power. While they provide a solid foundation, their performance may be limited when faced with complex patterns or nonlinear relationships.
@@
Machine learning methods have gained prominence in time-series forecasting due to rapid advancements in the field. Algorithms such as XGBoost [24], GBRT [25], Random Forests [26], and LightGBM [27] offer enhanced capabilities to handle nonlinear relationships and complex patterns. These methods demonstrate flexibility in handling different types and lengths of time-series data and generally provide superior forecasting accuracy compared to traditional statistical methods.
@@
Leveraging the representation learning capabilities of deep neural networks, Deep-Learning (DL) methods have shown promising results in time-series forecasting. DL treats time-series as sequences of vectors and utilizes architectures such as convolutional neural networks (CNNs) [28], recurrent neural networks (RNNs) [29], or Transformers [30] to capture temporal dependencies. For example, TCN [31] and DeepAR [32] implement CNNs or RNNs to model the temporal structure of the data. Transformer architectures, like REformer [33], Informer [34], Pyformer [35], FEDformer [15], Autoformer [13], Triformer [36], and PatchTST [37], have also been applied in time-series forecasting tasks, allowing for the capture of more complex temporal dynamics and significantly improving forecasting performance. In addition, while pursuing forecasting accuracy with complex models, MLP-based models such as N-HiTS [38], N-BEATS [39], and DLinear [40] employ a straightforward architecture with a relatively low number of parameters while achieving competitive performance. Recently, Mamba [41], a selective state space model, has also gained traction due to its ability to process dependencies in sequences while maintaining near-linear complexity. Some variants of Mamba [42, 43, 44] have been successfully applied to time-series forecasting.
2.2 Data-driven Numerical Weather Prediction
@@
Since 2022, there has been a growing interest in data-driven Numerical Weather Prediction (NWP) models within the AI and atmospheric science communities. The correlations between NWP and GSWF are as follows: 1) GSWF can be obtained by interpolating the forecast results of NWP models to specific latitudes and longitudes. Similarly, GSWF can also be used to bias-correct the forecast results of NWP models. These models, like, Pangu-Weather [9], GraphCast [8], FengWu [10], and FengWu-GHR [11], have shown the potential to outperform traditional physical-based NWP models in terms of forecast skill and operational efficiency. However, these models, operating at the mesh space ( e,g., the grid resolution of $0.25^{\circ}$ and $0.09^{\circ}$), are may not be the optimal solution for GSWF as discussed in Section 1. Prior to our work, some initial attempts, like Corrformer [12], have treated GSWF as an independent forecasting task, demonstrating promising results but remaining a great gap compared with the data-driven NWP models.
@@ 3 WEATHER-5K: Global Station Weather Dataset
3.1 Collection and Processing
Our dataset is collected using data sourced from the National Centers for Environmental Information (NCEI), specifically the Integrated Surface Database (ISD),  [^1] https://www.ncei.noaa.gov/products/land-based-station/integrated-surface-database a global repository of hourly and synoptic surface observations gathered from numerous original data sources. The ISD encompasses a wide range of meteorological parameters recorded by each weather station, including wind speed and direction, temperature, dew point, cloud data, sea level pressure, altimeter setting, station pressure, present weather, visibility, precipitation, snow depth, and various other elements. Although the ISD contains records from over $20,000$ stations spanning several decades, not all stations are suitable for machine learning applications. For instance, certain stations are no longer operational, many do not report data on an hourly basis, and numerous stations have missing values for critical weather elements. To create the WEATHER-5K dataset, a meticulous selection process was conducted to include only long-term, hourly reporting stations that are currently operational and provide essential observations such as temperature, dew point temperature, wind, and sea level pressure.
To obtain a high-quality dataset of weather stations, a series of post-processing steps were performed on the raw weather station data collected from 2014 to 2024. Initially, $10,701$ commonly operating stations were identified. The first step involved selecting stations that reported data every hour on the hour. However, many stations did not meet this criterion. To address this, a replacement method estimated missing hourly data points using the nearest available time points within a 30-minute window, significantly improving the distribution of valid hourly data.
@@ 3.1 Collection and Processing
Despite this improvement, a tiny portion of hourly data points remained missing due to the lack of observations within the 30-minute window. To fill these gaps, linear interpolation was employed using data from the 12 consecutive hours surrounding each missing point, ensuring the reliability of the interpolated data. Moreover, stations with more than 90% valid hourly data were chosen as the final candidates. To ensure the dataset’s completeness, ERA5 reanalysis data [45] was utilized to interpolate and fill any remaining gaps in the weather station data.
@@
As a result of these post-processing steps, a subset of $5,672$ weather stations worldwide was selected, spanning the period from 2014 to 2023, ensuring a recent and relevant time frame. This selection process focused on balancing the longevity of station operation, hourly data availability, and the inclusion of diverse weather variables. Consequently, our dataset provides a robust foundation for evaluating forecasting models in the context of GSWF.
@@ 3.2 Qualitative Analysis
Figure 2: The visualization of time-series data over a year. a), b) and c) represent the seasonal pattern for the same variables (i.e., temperature) at different regions (i.e., Southern Hemisphere, Northern Hemisphere, and Equatorial Regions). d), e), and f) represent the characteristics of different variables (i.e., wind direction, wind speed and sea-level temperature) at the same region (i.e., New York).
Regional disparities of stations. The WEATHER-5K dataset reveals regional disparities in the distribution of weather stations, which can significantly impact the learning and understanding of atmospheric dynamics in certain areas. As illustrated in Figure 1, some land regions have sparse data coverage compared to others. These disparities can be attributed to factors such as geographical characteristics, levels of economic development, and the strategic placement of weather stations. Note that the number of oceanic stations is also very limited due to the expensive cost of establishing stations at sea. Regions with limited station coverage may exhibit unique weather patterns and phenomena that are not adequately captured due to insufficient data. Addressing these disparities and expanding coverage in underrepresented areas is crucial for improving the accuracy and reliability of weather forecasting and analysis in those regions.
Data characteristics. Here we present a statistical analysis for different characteristics of different variables and the discrepancy of the same variable over different latitudes based on one year of data, By visualizing the four noteworthy observations: temperature, wind speed, wind direction, and sea level pressure, There are three primary findings: First, as shown in Figures 2 a), b), and c), temperature characteristics in the Northern and Southern Hemispheres are opposite during the same seasons. In the equatorial region, the variance for its seasonal pattern is generally lower compared to higher latitudes. Second, Figures 2 d) and e) indicate Wind Speed and Direction. They show that the wind speed and direction are non-stationary, characterized by intense fluctuations and a lack of clear patterns, making them challenging to predict. Last, As depicted in Figure 2 f), Sea Level Pressure shows a stable distribution with fewer fluctuations, making it relatively easier to predict.
@@
Exchange [46] Exchange 1 day 7,588 1 8 1990-2010 623KB
Electricity [47] Electricity 1 hour 26,304 321 1 2016-2019 92MB
ETTm2 [34] Electricity 15 mins 57,600 1 7 2016-2018 9.3MB
Traffic [13] Traffic 1 hour 17,544 862 1 2016-2018 131MB
LargeST-CA [48] Traffic 5 mins 525,888 8600 1 2017-2021 36.8GB
Solar [46] Weather 10 mins 52,560 137 1 2006 8.3MB
Wind [49] Weather 15 mins 48,673 1 7 2020-2021 2.7MB
Weather [13] Weather 10 mins 52,696 1 21 2020 7.0MB
Weather-Australia [14] Weather 1 day 1,332$\sim$65,981 3,010 4 unknown 202MB
GlobalTempWind [12] Weather 1 hour 17,544 3,850 2 2019-2020 1034MB
CMA_Wind [12] Weather 1 hour 17,520 34,040 1 2018-2019 N/A
WEATHER-5K (Ours) Weather 1 hour 87,648 5672 5 2014-2023 40.0GB
Comparison with existing datasets. Table 1 presents a comparison between the WEATHER-5K and other popular time-series datasets. The development of time-series datasets suffers from several limitations. 1) Small scale and Out-of-date: The mainstream time-series datasets [46, 47, 13] used for research purposes remain relatively small in scale. For instance, datasets related to electricity consumption or exchange rates are sparse or outdated, which limits the practical application of forecasting models. These two natures also hinder the exploration of more scientific challenges associated with time-series forecasting. 2) Lagging behind other fields: Compared to natural language processing and computer vision, the time-series forecasting domain has been slower in incorporating large-scale datasets. The use of extensive datasets (e.g., Common Crawl and LAION-5B [50]) in other fields have demonstrated unprecedented economic value and significantly advanced scientific discoveries. However, until recently, the first large-scale time-series dataset, LargeST [48], is only introduced. In contrast, WEATHER-5K addresses the limitation of small-scale datasets by providing a comprehensive and large-scale collection of weather station data. This abundance of data enables researchers to tackle more complex forecasting challenges and explore the dynamics of weather patterns on a global scale.

### v2

# How far are today’s time-series models from real-world weather forecasting applications?

@@ Abstract
The development of Time-Series Forecasting (TSF) techniques is often hindered by the lack of comprehensive datasets. This is particularly problematic for time-series weather forecasting, where commonly used datasets suffer from significant limitations such as small size, limited temporal coverage, and sparse spatial distribution. These constraints severely impede the optimization and evaluation of TSF models, resulting in benchmarks that are not representative of real-world applications, such as operational weather forecasting. In this work, we introduce the WEATHER-5K dataset, a comprehensive collection of observational weather data that better reflects real-world scenarios. As a result, it enables a better training of models and a more accurate assessment of the real-world forecasting capabilities of TSF models, pushing them closer to in-situ applications. Through extensive benchmarking against operational Numerical Weather Prediction (NWP) models, we provide researchers with a clear assessment of the gap between academic TSF models and real-world weather forecasting applications. This highlights the significant performance disparity between TSF and NWP models by analyzing performance across detailed weather variables, extreme weather event prediction, and model complexity comparison. Finally, we summarise the result into recommendations to the users and highlight potential areas required to facilitate further TSF research. The dataset and benchmark implementation are available at: https://github.com/taohan10200/WEATHER-5K.
@@ 1 Introduction
Global Station Weather Forecasting (GSWF) is essential for providing precise and timely weather information, with significant implications across various sectors such as aviation Gultepe et al. (2019), agriculture Ukhurebor et al. (2022), and energy Dehalwar et al. (2016). In particular, reliable forecasts are crucial for early warning systems, aiding in the preparation for natural disasters and extreme weather events, thereby safeguarding lives and property Wang et al. (2024b); Sillmann et al. (2017). At present, the most mature and mainstream approach to precise station-based weather forecasting is the use of Numerical Weather Prediction (NWP) models. These models can be further divided into physically-based NWP models Phillips (1956); Lynch (2008) and data-driven NWP models Lam et al. (2023); Bi et al. (2023); Chen et al. (2023); Han et al. (2024a). Despite their widespread deployment and reputation as the most accurate algorithms for station weather forecasting, NWP models are inherently computationally intensive. Specifically, the execution of these algorithms depends on vast resources for data assimilation and medium-range forecasting.
On the other hand, Time-Series Forecasting (TSF) methods, which offer a more economical solution, have also attracted attention from the weather forecasting community. Recently, many TSF methods have been developed and have demonstrated significant performance on small-scale weather station datasets. For instance, methods such as those presented in (Wu et al., 2021; Zhou et al., 2021; 2022; Zeng et al., 2023; Liu et al., 2024b) have achieved exceptionally low normalized errors (e.g., below 0.23 for temperature predictions). Despite their substantial success in research-oriented domains, the applicability of these methods for real-world weather forecasting remains under-explored. Specifically, most of these methods are trained and tested on single-station datasets Wu et al. (2021) or localized regions Godahewa et al. (2021), thereby limiting their generalizability to other stations. As depicted in the right panel of Figure 1, which tracks a winter storm forecast three days in advance, the TSF method exhibits a significant performance gap when tested on unseen stations compared to the real-world weather forecasts generated by NWP model.
Figure 1: A case study of winter storm forecasting using methods from two communities: NWP model and TSF model. It demonstrates a big performance discrepancy between them.
Hence, an emergent question arises: Are today’s time-series models ready for deployment in real-time weather forecasting? If not, how far are they from real-world applicability, and what can be done to improve them? To answer it, this work will deeply investigate three key aspects:
@@
1) Developing a Large-Scale, Comprehensive Global Station Weather Dataset. To fully unlock the potential of existing TSF methods and substantially improve their generalization ability. Specifically, we propose WEATHER-5K, a large-scale time-series dataset for sparse weather forecasting, comprising 5,672 weather stations worldwide. This dataset provides diverse weather conditions and 10 years of hourly data per station, enabling long-term pattern analysis and robust forecasting model development. Also, we conduct thorough data analysis to uncover trends, patterns, and correlations.
2) Conducting Extensive Assessments Between Existing TSF Algorithms and Real-World Applications Using NWP as Reference Models. This assessment will extend beyond those shallow attempts such as Wu et al. (2023), which only compared TSF methods to NWP models within a one-day lead time and did not use the most accurate real-time forecasting products (e.g., ERA5 prediction). Here, we implement a comprehensive set of widely recognized time-series forecasting methods across domains like traffic, weather, and electricity prediction. To this end, we establish a standardized evaluation framework to conduct extensive benchmark experiments on WEATHER-5K,
3) Fostering the Development of TSF for Real-World Weather Forecasting by Identifying future opportunities. We have at least four recommendations for future research:
  • The need to prioritize improving the long-term forecasting performance of TSF methods for their practical application in general weather forecasting.
  • Besides general weather forecasting, we find that existing TSF models still lag significantly in predicting extreme weather events, which require more focused attention.
  • Large-parameter models do not necessarily perform better. In our benchmarks. So developing efficient time-series modeling algorithms may yield greater benefits than merely increasing model complexity.
  • Leveraging TSF methods to integrate NWP forecasts as prior knowledge can effectively improve long-term forecasting capabilities.
Overall, we believe that our dataset and benchmarks will greatly facilitate future TSF research in weather forecasting. Our efforts will help researchers better evaluate and compare different algorithms, ultimately leading to improved forecasting techniques.
Figure 2: Flow diagram of the benchmark. a) Developing a downloading API to retrieve the raw ISD data and then do some pre-posting processing. b) Conducting rigorous quality control on selected stations to obtain a high-quality ISD subset. c) Using ERA5 to complete some missing data in the selected stations, which ensures 100% data completeness for training TSF models. d) Training and evaluating some main-stream TSF models with the basic metrics and a new proposed SEDI metric for extreme value evaluation.
@@ 2 Related Work
Time-series forecasting.
@@
Time-series forecasting involves many domains Godahewa et al. (2021); Chen et al. (2017); Bai et al. (2020); Qiu et al. (2024) and their methods have undergone three stages since its birth: statistical learning, machine learning, and deep learning.
@@
Statistical learning methods, including ARIMA Box & Pierce (1970), ETS Hyndman et al. (2008), StatsForecast Federico Garza (2022), VAR Godahewa et al. (2021), and Kalman Filter (KF) Harvey (1990), are among the early proposals and are widely utilized. These methods rely on historical data to predict future values and are based on the assumption that past observations hold predictive power. While they provide a solid foundation, their performance may be limited when faced with complex patterns or nonlinear relationships.
@@
Machine learning methods have gained prominence in time-series forecasting due to rapid advancements in the field. Algorithms such as XGBoost Chen & Guestrin (2016), GBRT Friedman (2001), Random Forests Breiman (2001), and LightGBM Ke et al. (2017) offer enhanced capabilities to handle nonlinear relationships and complex patterns. These methods demonstrate flexibility in handling different types and lengths of time-series data and generally provide superior forecasting accuracy compared to traditional statistical methods.
@@
Leveraging the representation learning capabilities of deep neural networks, Deep-Learning (DL) methods have shown promising results in time-series forecasting. DL treats time-series as sequences of vectors and utilizes architectures such as convolutional neural networks (CNNs) Lim & Zohren (2021), recurrent neural networks (RNNs) Hewamalage et al. (2021), or Transformers Wen et al. (2022) to capture temporal dependencies. For example, TCN Bai et al. (2018) and DeepAR Salinas et al. (2020) implement CNNs or RNNs to model the temporal structure of the data. Transformer architectures, like REformer Kitaev et al. (2020), Informer Zhou et al. (2021), Pyformer Liu et al. (2021), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Triformer Cirstea et al. (2022), and PatchTST Nie et al. (2022), have also been applied in time-series forecasting tasks, allowing for the capture of more complex temporal dynamics and significantly improving forecasting performance. In addition, while pursuing forecasting accuracy with complex models, MLP-based models such as N-HiTS Challu et al. (2023), N-BEATS Oreshkin et al. (2019), and DLinear Zeng et al. (2023) employ a straightforward architecture with a relatively low number of parameters while achieving competitive performance. Recently, Mamba Gu & Dao (2023), a selective state space model, has also gained traction due to its ability to process dependencies in sequences while maintaining near-linear complexity. Some variants of Mamba Wang et al. (2024c); Ahamed & Cheng (2024); Shi (2024) have been successfully applied to time-series forecasting.
Data-driven numerical weather prediction.
@@
Since 2022, there has been a growing interest in data-driven Numerical Weather Prediction (NWP) models within the AI and atmospheric science communities. The correlations between NWP and GSWF are as follows: 1) GSWF can be obtained by interpolating the forecast results of NWP models to specific latitudes and longitudes. Similarly, GSWF can also be used to bias-correct the forecast results of NWP models. These models, like, Pangu-Weather Bi et al. (2023), GraphCast Lam et al. (2023), FengWu Chen et al. (2023), and FengWu-GHR Han et al. (2024a), have shown the potential to outperform traditional physical-based NWP models in terms of forecast skill and operational efficiency. However, these models, operating at the mesh space ( e,g., the grid resolution of $0.25^{\circ}$ and $0.09^{\circ}$), are may not be the optimal solution for GSWF as discussed in Section 1. Prior to our work, some initial attempts, like Corrformer Wu et al. (2023), have treated GSWF as an independent forecasting task, demonstrating promising results but remaining a great gap compared with the data-driven NWP models.
@@ 3 WEATHER-5K: Global Station Weather Dataset
3.1 Collection and Pre-processing
Data source.
WEATHER-5K is derived from global near-surface in-situ observations, which is achieved in the publicly available Integrated Surface Database (ISD, ), leveraging data from high-quality observation networks. ISD is a global repository of hourly and synoptic surface observations, which encompasses a wide range of meteorological parameters. In this study, we study the common variables: the wind speed and direction, temperature, dew point, and sea level pressure.
Station selection.
ISD contains records from over $20,000$ stations spanning several decades, though certain stations are no longer operational, many do not report data on an hourly basis, and numerous stations have missing values for critical weather elements. To enhance data quality, a necessary meticulous selection process was conducted to include only long-term, hourly reporting stations that are currently operational and provide essential observations. As a result, there are $10,701$ operational stations from 2014 to 2024.
3.2 Quality Control
A high-quality dataset is crucial for TSF. As shown in Figure 2, WEATHER-5K has been subjected to rigorous post-processing and quality control to ensure the reliability of the dataset.
Data interpretation.
Due to the fact that each variable in the original ISD data is not directly recorded as floating-point data, but rather as strings, where each string encapsulates the numerical value, quality flag, reporting type, etc. of the variable. For example, a temperature field <+0130,1><+0130,1>< + 0130 , 1 > means the temperature value is 13.0 C, and it is at the first quality level. Therefore, the first step is to parse each variable according to the provided guidelines into a data format that is convenient for our understanding and storage.
Temporal alignment.
@@ 3.2 Quality Control
The aforementioned pre-processing has selected stations that report hourly data. However, some stations may not report at 1:00, 2:00, …, and are with a little time shift from the hour, such as 0:54, and 2:05. To address this, we merge the latest time as the hour to keep the temporal consistency. In implementation, we develop an algorithm that estimates missing hourly data points using the nearest available time points within a 30-minute window, significantly improving the distribution of valid hourly data. Despite this improvement, a tiny portion of hourly data points remained missing due to the lack of observations within the 30-minute window. To fill these gaps, linear interpolation was employed using data from the 12 consecutive hours surrounding each missing point, ensuring the reliability of the interpolated data.
Completeness filtering.
@@
To obtain a high-quality dataset of weather stations, only stations with more than 90% valid hourly data were chosen as the final candidates. As a result, $5,672$ weather stations worldwide are selected, spanning the period from 2014 to 2023, ensuring a recent and relevant time frame. This selection process focused on balancing the longevity of station operation, hourly data availability, and the inclusion of diverse weather variables.
Outlier detection.
The outlier detection process begins by meticulously examining the temporal dynamics of the dataset, focusing on identifying data points that lie far outside the expected range or exhibit unusual behaviors compared to the majority of observations. We use statistical techniques and machine learning algorithms to differentiate genuine anomalies from noise or data errors.
As the quality control process, only $5,672$ weather stations worldwide are selected, spanning the period from 2014 to 2023, ensuring a recent and relevant time frame. This selection process focused on balancing the longevity of station operation, hourly data availability, and the inclusion of diverse weather variables.
3.3 Post-Processing
There are still very few missing data points after quality control. This is mainly due to some stations experiencing long-term missing data (such as exceeding a full day). However, ensuring the completeness of data for all stations is particularly crucial for spatial modeling with the TSF algorithms. Therefore, we interpolate the ERA5 dataset Hersbach et al. (2018) to fill in the missing data points at the station locations. ERA5 is the most widely used high-quality reanalysis dataset globally, thus introducing only minor errors. However, since wind is a high-frequency parameter, the impact of using ERA5 for wind speed interpolation will be most significant.
To facilitate training on WEATHER-5K dataset, we further calculate the mean and variance of each variable over a decade to standardize the input data. Additionally, we observed a lack of evaluation metrics for extreme values in current TSF assessment methods, especially considering the crucial role of extreme values in weather forecasting. Therefore, we also computed percentiles at different thresholds ($90\%,95\%$, 98%, 99.5%) for each variable to propose metrics for extreme values.
@@ 3.4 Qualitative Analysis
Figure 3: a) and b) The visualization of time-series data over a year. c) The geographical distribution of the weather stations in WEATHER-5K. d) The error between the observations and the ERA5 dataset. e) The daily 2m temperature at station 57516099999 in Chongqing City from 1st June to 15th September, where filled areas represent the variance from the daily mean.
Distribution disparities of stations. The WEATHER-5K dataset reveals regional disparities in the distribution of weather stations, which can significantly impact the learning and understanding of atmospheric dynamics in certain areas. As illustrated in Figure 3 c), some land regions have sparse data coverage compared to others. These disparities can be attributed to factors such as geographical characteristics, levels of economic development, and the strategic placement of weather stations. Note that the number of oceanic stations is also very limited due to the expensive cost of establishing stations at sea. Regions with limited station coverage may exhibit unique weather patterns and phenomena that are not adequately captured due to insufficient data. Addressing these disparities and expanding coverage in underrepresented areas is crucial for improving the accuracy and reliability of weather forecasting and analysis in those regions.
Data patterns over one year.
Here we present a statistical analysis for different characteristics of different variables and the discrepancy of the same variable over different latitudes based on one year of data, By visualizing the two noteworthy observations: temperature, and wind speed. There are two primary patterns among different variables: First, as shown in Figure 3 a), temperature shows a seasonal pattern. Second, Figure 3 b) indicates wind speeds are non-stationary, characterized by intense fluctuations and a lack of clear patterns, making them challenging to predict.
Compare in-situ observations with ERA5.
ERA5 is a simulated dataset rather than being based on in-situ observations, which can limit its applicability in real-world scenarios. In contrast, WEATHER-5K dataset is derived from in-situ observations. To demonstrate their difference, we calculate the RMSE of ERA5 against in-situ observations (Jiao et al., 2021), as shown in Figure 3 d), which highlights an unacceptably high RMSE when comparing ERA5 data to real-world observations.
Heatwave analysis in China, 2022.
Figur 3 e) illustrates the daily temperature variations at Chongqing city in China (57516099999) from June to September in 2022. From this event, we can see that both the average and the maximum temperatures of WEATHER-5K are greater than those of in ERA5 in most of the heatwave days. This indicates that ERA5 consistently underestimated the diurnal temperature range at this station throughout the heatwave period.
@@
Exchange Exchange 1 day 7,588 1 8 1990-2010 623KB
Electricity Electricity 1 hour 26,304 321 1 2016-2019 92MB
ETTm2 Electricity 15 mins 57,600 1 7 2016-2018 9.3MB
Traffic Traffic 1 hour 17,544 862 1 2016-2018 131MB
LargeST-CA Traffic 5 mins 525,888 8600 1 2017-2021 36.8GB
Solar Weather 10 mins 52,560 137 1 2006 8.3MB
Wind Weather 15 mins 48,673 1 7 2020-2021 2.7MB
Weather Weather 10 mins 52,696 1 21 2020 7.0MB
Weather-Australia Weather 1 day 1,332$\sim$65,981 3,010 4 unknown 202MB
GlobalTempWind Weather 1 hour 17,544 3,850 2 2019-2020 1034MB
CMA_Wind Weather 1 hour 17,520 34,040 1 2018-2019 N/A
WEATHER-5K Weather 1 hour 87,648 5672 5 2014-2023 40.0GB
Comparison with existing datasets. Table 1 presents a comparison between the WEATHER-5K and other popular TSF datasets, which shows several limitations in previous datasets. 1) Small scale and out-of-date: The mainstream time-series datasets (Lai et al., 2018; Trindade, 2015; Wu et al., 2021) used for research purposes remain relatively small in scale. For instance, datasets related to electricity consumption or exchange rates are sparse or outdated, which limits the practical application of forecasting models. 2) Lagging behind other fields: The TSF community has been slower in incorporating large-scale datasets. The use of extensive datasets (e.g., Common Crawl and LAION-5B Schuhmann et al. (2022)) in other fields have demonstrated unprecedented economic value and significantly advanced scientific discoveries. However, until recently, the first large-scale time-series dataset, LargeST Liu et al. (2024a), is only introduced. The proposed WEATHER-5K will address the limitation of small-scale time series weather datasets. This abundance of data enables researchers to tackle more complex forecasting challenges. See Section C for more dataset analysis.